OpenStudy (anonymous):

Solve the system by substitution {3x+y-2x=22 {x+5y+z=4 {x=-3z

OpenStudy (nightwielder):

substitution means to plug in

OpenStudy (anonymous):

right

OpenStudy (nightwielder):

oops. x=-3z

OpenStudy (nightwielder):

plug in x=-3z (-3z)+5y+z=4 5y-2z=4 5y=2z+4 y=2/5z+4/5

OpenStudy (nightwielder):

plug x= and y= 3(3z)+(2/5z+4/5)-2z=22 9z-2z+2/5z+4/5=22

OpenStudy (anonymous):

I think I understand

OpenStudy (anonymous):

what else?

OpenStudy (nightwielder):

more coming

OpenStudy (nightwielder):



OpenStudy (nightwielder):